OpenStudy (anonymous):

I need help with this one sadly confused Simplify. (3m^2n^4)^4 81m^8n^16 12m^8n^16 81m^6n^8 12m^6n^8

OpenStudy (anonymous):

@ganeshie8 @texaschic101 @Preetha @tHe_FiZiCx99

ganeshie8 (ganeshie8):

hint : \[\huge (ab)^n = a^n.b^n\]

OpenStudy (anonymous):

the answer i got is A

OpenStudy (anonymous):

\[(3m^2n^4)^4 = 81m^{2 \times 4} n^{4 \times 4} = 81m^8n^{16}\] Got it

OpenStudy (anonymous):

so A is correct

OpenStudy (anonymous):

Yes